# The Population Resemblance Statistic: A Chi-Square Measure of Fit for Banking

### v3

@@ 3.4 Implementation guide
To implement the PRS framework in practice, a principled approach is essential, particularly in parameter selection. This section offers practical guidance for risk managers and analysts to ensure effective implementation across risk portfolios of varying sizes.
As informed by the asymptotic theory, the risk tolerance parameter $\delta$ depends inherently on the sample size $n$ and cannot be chosen arbitrarily. This parameter, $\delta$, which represents the maximum acceptable category-wise deviation, should be chosen inversely proportional to the sample size. Since the quantity $\{p_{0j}(1-p_{0j})/n\}^{1/2}$ corresponds to the standard error of a sample proportion, it provides a natural benchmark for determining meaningful deviations. We therefore propose defining $\delta$ as
@@
where $c>0$ is a relative scaling factor that adjusts the acceptable magnitude of shift in relation to this standard-error-like quantity. Analogous to the number of standard deviations from the model construction probabilities, $c$ provides a clear interpretation of how much deviation is considered acceptable. This formulation ensures that $\delta$ scales appropriately with sample size and category-specific uncertainty, serving as a measure for determining when the original model is no longer acceptable for use.
@@
For example, in a credit scoring model with probabilities $p_{0j}$ distributed across customer risk categories, $\delta$ ensures that no single category experiences significant shift – such as an increase in high-risk customers – large enough to compromise the confidence that the model can be expected to operate as intended. The PRS is calibrated to detect when one or more categories exhibit such levels of shift. By linking $\delta$ to sampling variability, this framework prevents substantive deviations across categories while maintaining a principled connection to sample size and category-specific uncertainty.
While the sections that follow restrict performance demonstration to equi-probable $\bm{p}_{0}$ for simplicity, the PRS procedure applies to arbitrary reference distributions. Note, however, that while the PRS statistic is asymptotically distribution-free, the decision boundaries determined according to (4) depend on $\bm{p}_{0}$. The recommendation in (4) is one proposed solution to appropriately adjust for sample size. Nevertheless, universal choices remain possible – one could always calculate $\delta$ using an equi-probable reference model, setting $\bm{p}_{0}=(1/B,\ldots,1/B)$ for the purpose of determining $\delta$, while remembering to use the true $\bm{p}_{0}$ when calculating the PRS statistic, if a predefined benchmark is preferred.
The multiplier $M>1$ must satisfy $M\delta\leq\min\limits_{j=1,\ldots,B}p_{0j}$ to ensure mathematical validity. The choice of $M$ depends on the context: smaller values are suitable for critical models requiring swift intervention, while larger values allow for more flexibility in settings where greater shifts can be tolerated before corrective action is needed. Setting excessively large $M$ (relative to $c$) can distort the framework by shifting the critical values too far apart, undermining the reliability of monitoring decisions.
@@
Practically, the PRS is calibrated so that shifts statistically equivalent to $n\delta$ cases per risk category signal a transition from acceptable to partial discrepancy. Similarly, shifts statistically equivalent to $nM\delta$ cases per risk category mark the shift from partial to full discrepancy. While $n\delta$ and $nM\delta$ provide easy reference points, actual deviations vary across categories – some shifts less, others more – necessitating a statistical procedure to assess overall resemblance.
@@
The sensitivity parameters $\alpha_{1}$ and $\alpha_{2}$ should be chosen based on the organization’s desired balance between stickiness and responsiveness, ensuring consistent control over decision boundaries across monitoring regions – lower values prioritize model continuity while higher values lead to increased sensitivity to population shifts. Additionally, while $M$ and $c$ can be tailored to each portfolio being monitored, a more robust approach – one that mitigates the risk of cherry-picking results – is to fix these parameters across all portfolios, as $\delta$ already accounts for sample size considerations. This approach simplifies implementation while maintaining a systematic and objective monitoring framework.
@@ 4 PRS method competency: A simulation study
For each scenario, the reference probability distribution was defined as $\bm{p}_{0}=(1/B,\dots,1/B)$, representing $B$ equally likely categories. A deviation of magnitude $\delta_{v}$ was introduced to define perturbed probability distributions $\bm{p}_{v}$, representing the current model, given by
$p_{v,j}=\begin{cases}1/B-\delta_{v},&j\leq B/2,\\ 1/B+\delta_{v},&j\geq B/2+1,\end{cases}$
where, when $B$ is odd, the central category remains unchanged at $1/B$. Under this setting, the PRS tolerance level defined in (4) becomes $\delta=cB^{-1}\sqrt{(B-1)/n}$ with specified scaling constant $c$. The boundary for unacceptable deviation is set at $M\delta$ for specified $M>1$. Using the decision-making framework of Definition 2, we establish the PRS critical values $\tau_{1}$ and $\tau_{2}$, delineating the decision regions $\mathcal{R}_{1}$, $\mathcal{R}_{2}$ and $\mathcal{R}_{3}$.
We evaluated the PRS classification probabilities across 30 values of $\delta_{v}$, ranging from no deviation, $\delta_{v}=0$, to an extreme deviation, $\delta_{v}=(3M+2)\delta$. This range captures shifts from within the acceptable range, where the population remains resemblant to the model, to larger deviations that are fully discrepant, indicating a substantial shift and necessitating model reconstruction.
@@
Simulations were conducted across a range of values for $(n,B)$, varying $0.5\leq c\leq 1$, $21.2<M<$ 2, and $0.01\leq\alpha_{1},\alpha_{2}\leq 0.2$. For illustrative purposes, results are presented for $(n,B)=(50,5)$ under two scenarios, firstly $(c,M)=(0.7,2)$ and $(\alpha_{1},\alpha_{2})=(0.1,0.05)$ – see Figure 5 – and, secondly, $(c,M)=(1,1.6)$ and $(\alpha_{1},\alpha_{2})=(0.1,0.2)$ – see Figure 6. This is to illustrate the versatility of the method across parameter choices. Finally, for $(n,B)=(10,000,20)$ with, $(c,M)=(0.7,2)$ and $(\alpha_{1},\alpha_{2})=(0.05,0.1)$, we show the results in Figure 7. The empirical classification probabilities were plotted against $\delta_{v}$ to illustrate classification behavior across the full range of model deviations.
Figure 5: Empirical classification probabilities for
$(n,B)=(50,5)$ with $(\alpha_{1},\alpha_{2})=(0.1,0.05)$.
Figure 6: Empirical classification probabilities for
$(n,B)=(50,5)$ with $(\alpha_{1},\alpha_{2})=(0.1,0.2)$
Figure 7: Empirical classification probabilities for
$(n,B)=(10,000,20)$ with $(\alpha_{1},\alpha_{2})=(0.05,0.1)$.
The results demonstrate the effectiveness of the PRS method in identifying deviations that warrant model reconstruction. When $\delta_{v}=\delta$, the probability of PRS falling above the upper critical value $\tau_{2}$, $P(\mathcal{R}_{3})$, closely matches the expected value of $\alpha_{1}=0.05$. Similarly, under a deviation of size $\delta_{v}=M\delta$, the probability of falling below $\tau_{1}$, $P(\mathcal{R}_{1})$, remains near $\alpha_{2}=0.1$. These findings confirm the validity of the PRS procedure and its effectiveness in classifying multinomial samples under structured deviations from a reference model.
@@ 5 Banking applications and performance analysis
This section presents an empirical validation of the PRS based on a set of anonymized credit risk models using data from a large South African financial service provider. The dataset, spanning retail and corporate portfolios, was selected to evaluate the PRS methodology across operationally relevant scenarios. Portfolio sizes have been rounded as part of the anonymization effort. The examples encompass multiple portfolio configurations varying in sample size ($n$) and number of risk categories ($B$) where using equi-probable risk categories, ${p_{0i}}=1/B,\ i=1,...,B$, was deemed appropriate. Results are reported in Tables 3 through 5; the commonly used red-amber-green ($\mathbbm{rag}$) status as used in banks was assigned ($\mathbbm{r}$ indicating membership to $\mathcal{R}_{3}$, $\mathbbm{a}$ to $\mathcal{R}_{2}$, and $\mathbbm{g}$ to $\mathcal{R}_{1}$). The designs considered are $(n,B)\in\{(50,5),\,(500,10),\,(2,000,10),\,(10,000,20)\}$.
@@
We follow the procedure of Figure 4 to implement the PRS, using the recommended $\delta$ from (4) and choosing sensitivity parameters $\alpha_{1}=0.1$ and $\alpha_{2}=0.05$. These choices reflect risk aversion for an erroneous decision: $\alpha_{1}$ represents a $10\%$ probability of an $\mathcal{R}_{3}$ (fully discrepant) classification under $\delta$-resemblance, while $\alpha_{2}$ represents a $5\%$ probability of an $\mathcal{R}_{1}$ (acceptable) classification under $M\delta$-resemblance.
For these examples, we set $c=0.7$ and $M=2$, reflecting an institutionally acceptable level of risk tolerance. Conceptually, shifts statistically equivalent to $n\delta$ cases across all risk categories mark the transition from acceptable to partial discrepancy, while shifts statistically equivalent to $nM\delta$ cases across all risk categories mark the transition to full discrepancy. For $(n,B)=(50,5)$ and $(n,B)=(10,000,20)$, these values are $1.98\approx 2$ and $15.26\approx 15$ cases per category for the transition to partial discrepancy, and $3.96\approx 4$ and $30.51\approx 31$ cases per category for full discrepancy.
@@
The critical values $\tau_{1}$ and $\tau_{2}$ delineating the regions $\mathcal{R}_{1}\ (\mathbbm{g}reen),\ \mathcal{R}_{2}\ (\mathbbm{a}mber)$ and $\mathcal{R}_{3}\ (\mathbbm{r}ed)$ are tabulated in Table 2 along with the corresponding values of $\delta$. In all cases, $\delta$ (and $M\delta$) is smaller than $1/B$ ($=\min\limits_{j=1,\ldots,B}p_{0j}$), which is practically sensible.
We compare the outcomes of using the PRS with the widely used PSI, employing both the Lewis constants and the critical values proposed by Yurdakul and Naranjo, (2020), denoted by “L” and “YN”, respectively. Of course, a direct comparison of the PRS with the PSI (using YN critical values) is not entirely fair because the latter is governed by the specification of two type I errors under a null hypothesis of exact equality. Moreover, the PSI(YN) is known to have weak power at small sample sizes (Table 4 in Yurdakul and Naranjo, (2020)), a limitation of PSI(YN) and perhaps explainable by the instability of the normalized PSI in small samples observed in Figures 2 and 2. The YN-normed critical values are $\tau=2n^{-1}F^{-1}_{B-1}(1-\alpha)$ (in the notation of Yurdakul and Naranjo, (2020) and assuming $m=n$), where $F_{\nu}^{-1}$ denotes the chi-square quantile function with $\nu$ degrees of freedom. We choose upper and lower significance levels of $1\%$ and $10\%$, respectively (if $p<1\%$ then $\mathbbm{r}$, if $p>10\%$ then $\mathbbm{g}$, otherwise $\mathbbm{a}$).
@@
A further comparison of the PRS is made with the discrete Kolmogorov-Smirnov (KS) test, although also not entirely fair given a null hypothesis of exact equality. Given the discrete nature of the KS statistic, particularly in small samples, it is not feasible to match critical values at specified significance levels and we rather report the p-value, $p(\mathrm{KS})$ from the dgof [^2] The syntax used is ks.test(X, ecdf(1:K), exact=F, simulate.p.value=T, B=10000) where X is a numeric vector containing the repeated category number based on $\hat{p}$ and $B$ therein is the number of simulations and K the number of categories. package in R, version 1.5.1 (2024/10/09), based on Arnold and Emerson, (2011). We also used the upper and lower significance levels of $1\%$ and $10\%$, respectively. Keep in mind that the significance levels $\alpha$ used for PSI(YN) and KS do not hold the same interpretation as $\alpha_{1}$ and $\alpha_{2}$ for the PRS.
Table 2: Critical values for the PRS, specifying $c=0.7$ and $M=2$, and $\alpha_{1}=10\%$ and $\alpha_{2}=5\%$, for $\delta$ from (4).
@@
$(50,5)$ 0.056569 0.07441 0.25722
$(500,10)$ 0.013416 0.03063 0.04890
(2 000,10)200010(2\ 000,10)( 2 000 , 10 ) 0.006708 0.00766 0.01222
(10 000,20)1000020(10\ 000,20)( 10 000 , 20 ) 0.002179 0.00394 0.00439
Table 3: Population resemblance comparison using the current population of size $n=50$ and observed category sample sizes $n_{i}$, $i=1,...,B$ for $B=5$. For all categories, $n_{0i}=10$.
${\mathbf{n_{i}}}$ PSI (L,YN) PRS p(KS)
@@
$t_{6}$ (2, 5, 13, 14, 16) 0.426 ($\mathbbm{r,a}$) 0.300 ($\mathbbm{r}$) $<0.001$ ($\mathbbm{r}$)
Table 4: Population resemblance comparison using the current population of size $n=500$ and observed category sample sizes $n_{i}$, $i=1,...,B$, $B=10$ and $n_{0i}=50$.
${\mathbf{n_{i}}}$ PSI (L,YN) PRS p(KS)
$t_{1}$Centerstack(35, 40, 45, 45, 47, 50, 55, 58, 60, 65) 0.032 ($\mathbbm{g,g}$) 0.032 ($\mathbbm{a}$) 0.002 ($\mathbbm{r}$)
$t_{2}$Centerstack(40, 45, 45, 45, 47, 48, 55, 55, 60, 60) 0.017 ($\mathbbm{g,g}$) 0.018 ($\mathbbm{g}$) 0.024 ($\mathbbm{a}$)
$t_{3}$Centerstack(35, 36, 42, 43, 44, 44, 60, 60, 61, 75) 0.060 ($\mathbbm{g,a}$) 0.062 ($\mathbbm{r}$) $<0.001$ ($\mathbbm{r}$)
$t_{4}$Centerstack(20, 35, 35, 40, 40, 62, 65, 65, 65, 73) 0.131 ($\mathbbm{a,r}$) 0.116 ($\mathbbm{r}$) $<0.001$ ($\mathbbm{r}$)
Table 5: Population resemblance comparison using the current population of size $n=2,000$ and observed category sample sizes $n_{i}$, $i=1,...,B$, $B=10$ and $n_{0i}=200$.
${\mathbf{n_{i}}}$ PSI (L,YN) PRS p(KS)
$t_{1}$Centerstack(160, 170, 180, 180, 190, 200, 210, 220, 240, 250) 0.020 ($\mathbbm{g,a}$) 0.020 ($\mathbbm{r}$) $<0.001$ ($\mathbbm{r}$)
$t_{2}$Centerstack(180, 180, 184, 190, 194, 200, 200, 210, 222, 240) 0.008 ($\mathbbm{g,g}$) 0.008 ($\mathbbm{a}$) 0.004 ($\mathbbm{r}$)
$t_{3}$Centerstack(180, 180, 190, 194, 200, 200, 204, 210, 220, 222) 0.005 ($\mathbbm{g,g}$) 0.005 ($\mathbbm{g}$) 0.035 ($\mathbbm{a}$)
$t_{4}$Centerstack(160, 170, 170, 178, 180, 210, 210, 220, 242, 260) 0.025 ($\mathbbm{g,r}$) 0.026 ($\mathbbm{r}$) $<0.001$ ($\mathbbm{r}$)
Table 6: Population resemblance comparison using the current population of size $n=10,000$ and observed category sample sizes $n_{i}$, $i=1,...,B$, $B=20$ and $n_{0i}=500$.
@@
$t_{1}$Centerstack(425, 455, 480, 480, 480, 480, 485, 491, 495, 495,
500, 502, 502, 502, 502, 520, 540, 546, 550, 570) 0.0042 ($\mathbbm{g,g}$) 0.0042 ($\mathbbm{a}$)
[…]
545, 545, 550, 550, 600, 620, 650, 650, 650, 650) 0.1060 ($\mathbbm{a,r}$) 0.0769 ($\mathbbm{r}$)
[…]
501, 502, 502, 510, 510, 520, 520, 530, 540, 550) 0.0025 ($\mathbbm{g,g}$) 0.0025 ($\mathbbm{g}$)
[…]
490, 490, 525, 525, 555, 555, 585, 585, 600, 600) 0.0139 ($\mathbbm{g,r}$) 0.0142 ($\mathbbm{r}$)
[…]
525, 525, 545, 545, 550, 550, 555, 555, 600, 600) 0.0153 ($\mathbbm{g,r}$) 0.0150 ($\mathbbm{r}$)
[…]
490, 490, 510, 510, 520, 520, 550, 550, 550, 575) 0.0045 ($\mathbbm{g,g}$) 0.0045 ($\mathbbm{r}$)
@@
When comparing the PSI(L) with the PRS, the results are commensurate with the conclusions from Table 1: the PRS procedure less frequently indicates “full discrepancy” in small samples ($n=50$), while doing so more frequently in larger samples ($n\geq 500$). The same conclusion holds when comparing the PSI(L) with the PSI(YN). Recall that the PSI(L) has an inflated probability of indicating a shift when none has occurred in small samples and, conversely, a too-close-to-zero probability of indicating “full discrepancy” in large samples. Notably, in the cases where $n$ exceeds $500$, PSI(L) seems insensitive to discrepancy (partial and full). Observe in all cases considered here, that the PRS indicates discrepancy ($\mathbbm{r}$, or $\mathbbm{a}$) more frequently than the PSI(YN).

### v4

@@ 3.4 Implementation guide
Implementing the PRS framework in practice requires a principled approach, especially in parameter selection. This section offers practical guidance for risk managers and analysts to ensure valid and effective implementation of the PRS across risk portfolios of varying sizes. Central is the tolerable shift parameter $\delta$, which defines the maximum acceptable category-wise deviation. This value is not freely specifiable: it must respect theoretical constraints imposed by the sample size $n$ and the number of categories $B$, ensuring the validity of the underlying non-central chi-square approximation.
Motivated by the asymptotic theory, we recommend defining $\delta$ as
@@
where $c>0$ is a scaling factor that adjusts the acceptable magnitude of shift in relation to sampling variability. Since the quantity $\{p_{0j}(1-p_{0j})/n\}^{1/2}$ corresponds to the standard error of a sample proportion, it provides a natural benchmark for defining what constitutes a meaningful deviation. In this setup, $c$ functions like a tolerance multiplier — comparable to the number of standard errors deemed acceptable — supporting consistent application across portfolios with different sizes and risk profiles.
@@
For example, in a credit scoring context, $\delta$ limits category-specific shifts (e.g., a rise in high-risk borrowers) that would undermine confidence in the model’s continued validity. The PRS is calibrated to detect when one or more categories exhibit such levels of shift. Tying $\delta$ to sampling variability guards against substantive shifts across categories while maintaining the connection to sample size and category-specific uncertainty.
While examples in later sections assume an equi-probable $\bm{p}_{0}$ for simplicity, the PRS applies to arbitrary reference distributions. Importantly, although the PRS statistic is asymptotically distribution-free, the decision boundaries defined using $\delta$ in (5) do depend on $\bm{p}_{0}$. The recommendation in (5) is but one possible solution to appropriately adjust for sample size. However, this flexibility is moderated by the mathematical constraint that $M\delta\leq\min_{j}p_{0j}$. This ensures that perturbed category probabilities remain valid.
As a result, for highly imbalanced reference distributions (i.e., when some $p_{0j}$ are small), the allowable range of $\delta$ may become narrow, limiting the magnitude of detectable shifts. The parameter $M>1$, which controls the threshold between partial and full discrepancy, should be chosen accordingly. Smaller values of $M$ are appropriate in settings that demand rapid response, while larger values allow for more tolerance before action is triggered. Care should be taken, however: setting $M$ excessively high (relative to $c$) may push critical values too far apart, reducing the reliability of monitoring decisions.
@@
Practically, the PRS is calibrated so that shifts statistically equivalent to $n\delta$ cases per risk category signal a transition from acceptable to partial discrepancy. Similarly, shifts statistically equivalent to $nM\delta$ cases per risk category mark the shift from partial to full discrepancy. While $n\delta$ and $nM\delta$ provide easy reference points, actual deviations vary across categories — some shifts less, others more — necessitating a statistical procedure to assess overall resemblance.
A universal alternative is to compute $\delta$ using a uniform reference, i.e., setting $\bm{p}_{0}=(1/B,\ldots,1/B)$ solely for the purpose of defining $\delta$, while still using the actual $\bm{p}_{0}$ when evaluating the PRS statistic. This approach can be useful when a fixed benchmark is desired across portfolios, but may distort the interpretation of shift sizes, especially when some $p_{0,j}$ are small. In such cases, common in portfolios with small sample sizes or high concentration risk, the uniform threshold can misrepresent meaningful deviations.
@@
The sensitivity parameters $\alpha_{1}$ and $\alpha_{2}$ should be chosen based on the organization’s desired balance between under- and overreaction, ensuring consistent control over decision boundaries across monitoring regions: lower values prioritize model continuity, while higher values lead to increased sensitivity to population shifts. Practitioners should caution against choosing excessively large values for $\alpha_{1}$ and $\alpha_{2}$ which may cause the intermediate monitoring region $\mathcal{R}_{2}$ to shrink or disappear entirely if the decision boundaries overlap. Our findings suggest that values in the range $[0.01,0.2]$ generally yield stable results, though a trade-off may be necessary if greater sensitivity is required in one direction or the other. Furthermore, setting excessively large $M$ can distort the framework by shifting the critical values too far apart, undermining the reliability of monitoring decisions. Additionally, while $M$ and $c$ can be tailored to each portfolio being monitored, a more robust approach — one that mitigates the risk of cherry-picking results — is to fix these parameters across all portfolios, as $\delta$ already accounts for sample size considerations. This approach simplifies implementation while maintaining a systematic and objective monitoring framework.
@@ 4 PRS method competency: A simulation study
For each scenario, the reference probability distribution was defined as $\bm{p}_{0}=(1/B,\dots,1/B)$, representing $B$ equally likely categories. A deviation of magnitude $\delta_{v}$ was introduced to define perturbed probability distributions $\bm{p}_{v}$. Specifically, the first half of the categories have reduced probabilities $p_{v,j}=1/B-\delta_{v}$ for $j=1,\ldots,\lfloor B/2\rfloor$, while the remaining categories have increased probabilities $p_{v,j}=1/B+\delta_{v}$ for $j=\lceil B/2\rceil+1,\ldots,B$. When $B$ is odd, the central category remains unchanged at $1/B$. Under this setting, the PRS tolerance level defined in (5) becomes $\delta=cB^{-1}\sqrt{(B-1)/n}$ with specified scaling constant $c$. The boundary for unacceptable deviation is set at $M\delta$ for specified $M>1$. Using the decision-making framework of Definition 2, we establish the PRS critical values $(\tau_{1},\tau_{2}$), delineating the decision regions $(\mathcal{R}_{1},\mathcal{R}_{2},\mathcal{R}_{3})$.
We evaluated the PRS classification probabilities across 30 values of $\delta_{v}$, ranging from no deviation, $\delta_{v}=0$, to an extreme deviation, $\delta_{v}=(3M+2)\delta$. This range captures shifts from within the acceptable range, where the population remains resemblant to the model, to larger deviations that are fully discrepant.
@@
Simulations were conducted across a range of values for $(n,B)$, varying $0.5\leq c\leq 1$, $1.2<M<2$, and $0.01\leq\alpha_{1},\alpha_{2}\leq 0.2$. For illustrative purposes, results are presented for $(n,B)=(50,5)$ under two scenarios, firstly $(c,M)=(0.7,2)$ and $(\alpha_{1},\alpha_{2})=(0.05,0.1)$ — see Figure 5 — and, secondly, $(c,M)=(1,1.6)$ and $(\alpha_{1},\alpha_{2})=(0.1,0.2)$ — see Figure 6. This is to illustrate the versatility of the method across parameter choices. Finally, for (n,B)=(10,000,20)(n,B)=(10{,}000,20) with, $(c,M)=(0.7,2)$ and $(\alpha_{1},\alpha_{2})=(0.05,0.1)$, we show the results in Figure 7. In each case, empirical classification probabilities are plotted against $\delta_{v}$ to visualize how classification behavior evolves across the deviation spectrum.
Figure 5: Empirical probabilities for $(n,B)=(50,5)$ with $(\alpha_{1},\alpha_{2})=(0.05,0.1)$.
Figure 6: Empirical probabilities for $(n,B)=(50,5)$ with $(\alpha_{1},\alpha_{2})=(0.1,0.2)$.
Figure 7: Empirical probabilities for (n,B)=(10,000,20)(n,B)=(10{,}000,20) with $(\alpha_{1},\alpha_{2})=(0.05,0.1)$.
The results demonstrate that the PRS method effectively identifies when deviations become severe enough to warrant model reconstruction. When $\delta_{v}=\delta$, the probability of PRS falling above the upper critical value $\tau_{2}$, $P(\mathcal{R}_{3})$, closely matches the expected value of $\alpha_{1}$. Similarly, under a deviation of size $\delta_{v}=M\delta$, the probability of falling below $\tau_{1}$, $P(\mathcal{R}_{1})$, remains near $\alpha_{2}$. These findings support the validity of the PRS framework and its reliability in classifying structured deviations from a multinomial reference model.
@@ 5 Banking applications and performance analysis
This section presents an empirical validation of the PRS based on a set of anonymized credit risk models using data from a large South African financial service provider. The dataset, spanning retail and corporate portfolios, was selected to evaluate the PRS methodology across operationally relevant scenarios. Portfolio sizes have been rounded as part of the data anonymization. The examples encompass multiple portfolio configurations varying in sample size ($n$) and number of risk categories ($B$) where using equi-probable risk categories, ${p_{0i}}=1/B,\ i=1,...,B$, was deemed appropriate. Results are reported in Tables 3 and 4. A red-amber-green ($\mathbbm{RAG}$) status was assigned using the decision regions in Definition 2: red indicates full discrepancy, amber partial discrepancy, and green no significant shift. The designs considered are $(n,B)\in\{(50,5),\,(500,10),\,(2{,}000,10),\,(10{,}000,20)\}$.
Note, category counts $n_{i}$ are tabulated rather than proportions $\hat{p}_{i}$ to enhance readability. This does not imply that $n$ remains fixed over time, nor that counts are the key statistic; the monitoring objective remains the evolution of $\hat{\bm{p}}$ relative to $\bm{p}_{0}$.
@@
We follow the procedure of Figure 4 to implement the PRS, using the recommended $\delta$ from (5) and choosing sensitivity parameters $\alpha_{1}=0.05$ and $\alpha_{2}=0.1$. These choices reflect risk aversion for an erroneous decision: $\alpha_{1}$ represents a $5\%$ probability of a red classification under $\delta$-resemblance, while $\alpha_{2}$ represents a $90\%$ probability of an amber or red classification under $M\delta$-resemblance. We used $c=0.7$ and $M=2$ for these applications.
@@
Conceptually, shifts statistically equivalent to $n\delta$ cases across all risk categories mark the transition from green to amber, while shifts statistically equivalent to $nM\delta$ cases across all risk categories mark the transition to red. For $(n,B)=(50,5)$ and (n,B)=(10,000,20)(n,B)=(10{,}000,20), these values are $1.98\approx 2$ and $15.26\approx 15$ cases per category for the transition to amber, and $3.96\approx 4$ and $30.51\approx 31$ cases per category to red. The critical values $\tau_{1}$ and $\tau_{2}$ delineating the regions green, amber and red are tabulated in Table 2 along with the corresponding values of $\delta$. In all cases, $\delta$ (and $M\delta$) is smaller than $1/B$ ($\min\limits_{j=1,\ldots,B}p_{0j}$), which is practically sensible.
We compare the PRS results with the widely used PSI, employing both the Lewis constants and the critical values proposed by [yurdakul2020statistical] (denoted by “L” and “YN”, respectively). We also include the discrete Kolmogorov-Smirnov (KS) test for further comparison. Of course, a direct comparison of the PRS with the PSI(YN) or the discrete KS is not entirely equivalent — these are governed by the specification of two type I errors under a point null hypothesis of exact equality, in contrast to the nested composite structure of our proposed procedure. Nonetheless, we choose upper and lower significance levels of $1\%$ and $10\%$: $p<1\%$ yields red, $p>10\%$ green, and intermediate values amber.
[…]
Moreover, it is known that PSI(YN) has low power at small sample sizes [yurdakul2020statistical, Table 4], possibly due to instability of the normalized PSI, as also seen in Figures 2 and 2. The YN-normed critical values are $\tau=2n^{-1}F^{-1}_{B-1}(1-\alpha)$ (in the notation of [yurdakul2020statistical] and assuming $m=n$), where $F_{\nu}^{-1}$ denotes the chi-square quantile function with $\nu$ degrees of freedom.
@@
Given the discrete nature of the KS statistic, particularly in small samples, it is not feasible to match critical values at specified significance levels and we rather report the p-value, $p(\mathrm{KS})$ from the dgof [^2] The syntax used is ks.test(X, ecdf(1:K), exact=F, simulate.p.value=T, B=10000) where X is a numeric vector containing the repeated category number based on $\hat{p}$ and $B$ therein is the number of simulations and K the number of categories. package in R, version 1.5.1 (2024/10/09), based on [RDGOF].
Table 2: Critical values for the PRS, specifying $(c,M)=(0.7,2)$ and $(\alpha_{1},\alpha_{2})=(5\%,10\%)$, for $\delta$ from (5).
@@
$(50,5)$ 0.039598 0.07441 0.25722
$(500,10)$ 0.009391 0.03063 0.04890
(2 000,10)(2\ 000,10) 0.004696 0.00766 0.01222
(10 000,20)(10\ 000,20) 0.001526 0.00394 0.00439
Table 3: Population resemblance comparison using the current population of size $n$ and observed category sample sizes $n_{i}$, i=1,,Bi=1,...,B.
3A $\mathbf{n_{i}}$} with $n_{0i}=10$, $(n,B)=(50,5)$ PSI (L,YN) PRS p(KS)
@@
$t_{6}$ (2, 5, 13, 14, 16) 0.426 ($\mathbbm{R,A}$) 0.300 ($\mathbbm{R}$) <0.001<0.001 ($\mathbbm{R}$)
3B $\mathbf{n_{i}}$} with $n_{0i}=50$, $(n,B)=(500,10)$ PSI (L,YN) PRS p(KS)
$t_{1}$ (35, 40, 45, 45, 47, 50, 55, 58, 60, 65) 0.032 ($\mathbbm{G,G}$) 0.032 ($\mathbbm{A}$) 0.002 ($\mathbbm{R}$)
$t_{2}$ (40, 45, 45, 45, 47, 48, 55, 55, 60, 60) 0.017 ($\mathbbm{G,G}$) 0.018 ($\mathbbm{G}$) 0.024 ($\mathbbm{A}$)
$t_{3}$ (35, 36, 42, 43, 44, 44, 60, 60, 61, 75) 0.060 ($\mathbbm{G,A}$) 0.062 ($\mathbbm{R}$) <0.001<0.001 ($\mathbbm{R}$)
$t_{4}$ (20, 35, 35, 40, 40, 62, 65, 65, 65, 73) 0.131 ($\mathbbm{A,R}$) 0.116 ($\mathbbm{R}$) <0.001<0.001 ($\mathbbm{R}$)
3C $\mathbf{n_{i}}$} with $n_{0i}=200$, (n,B)=(2,000,10)(n,B)=(2{,}000,10) PSI (L,YN) PRS p(KS)
$t_{1}$ (160, 170, 180, 180, 190, 200, 210, 220, 240, 250) 0.020 ($\mathbbm{G,A}$) 0.020 ($\mathbbm{R}$) <0.001<0.001 ($\mathbbm{R}$)
$t_{2}$ (180, 180, 184, 190, 194, 200, 200, 210, 222, 240) 0.008 ($\mathbbm{G,G}$) 0.008 ($\mathbbm{A}$) 0.004 ($\mathbbm{R}$)
$t_{3}$ (180, 180, 190, 194, 200, 200, 204, 210, 220, 222) 0.005 ($\mathbbm{G,G}$) 0.005 ($\mathbbm{G}$) 0.035 ($\mathbbm{A}$)
$t_{4}$ (160, 170, 170, 178, 180, 210, 210, 220, 242, 260) 0.025 ($\mathbbm{G,R}$) 0.026 ($\mathbbm{R}$) <0.001<0.001 ($\mathbbm{R}$)
Table 4: Population resemblance comparison using the current population of size n=10,000n=10{,}000 and observed category sample sizes $n_{i}$, i=1,,Bi=1,...,B, $B=20$ and $n_{0i}=500$.
@@
$t_{1}$            500, 502, 502, 502, 502, 520, 540, 546, 550, 570)(425, 455, 480, 480, 480, 480, 485, 491, 495, 495, 0.0042 ($\mathbbm{G,G}$) 0.0042 ($\mathbbm{A}$)
$t_{2}$            545, 545, 550, 550, 600, 620, 650, 650, 650, 650)(150, 170, 400, 400, 450, 450, 460, 460, 525, 525, 0.1060 ($\mathbbm{A,R}$) 0.0769 ($\mathbbm{R}$)
$t_{3}$            501, 502, 502, 510, 510, 520, 520, 530, 540, 550)(445, 455, 480, 480, 485, 485, 490, 495, 500, 500, 0.0025 ($\mathbbm{G,G}$) 0.0025 ($\mathbbm{G}$)
$t_{4}$            490, 490, 525, 525, 555, 555, 585, 585, 600, 600)(425, 425, 440, 440, 445, 445, 460, 460, 475, 475, 0.0139 ($\mathbbm{G,R}$) 0.0142 ($\mathbbm{R}$)
$t_{5}$            525, 525, 545, 545, 550, 550, 555, 555, 600, 600)(390, 390, 450, 450, 450, 450, 460, 460, 475, 475, 0.0153 ($\mathbbm{G,R}$) 0.0150 ($\mathbbm{R}$)
$t_{6}$            490, 490, 510, 510, 520, 520, 550, 550, 550, 572)(440, 465, 465, 475, 475, 480, 480, 485, 485, 488, 0.0045 ($\mathbbm{G,G}$) 0.0045 ($\mathbbm{R}$)
@@
When comparing the PSI(L) with the PRS, the results are commensurate with the conclusions from Table 1: the PRS procedure less frequently indicates “full discrepancy” (red) in small samples ($n=50$), while doing so more frequently in larger samples ($n\geq 500$). The same conclusion holds when comparing the PSI(L) with the PSI(YN). Recall that the PSI(L) has an inflated probability of indicating a shift when none has occurred in small samples and, conversely, a too-close-to-zero probability of indicating “full discrepancy” (red) in large samples. Notably, in the cases where $n$ exceeds $500$, PSI(L) seems insensitive to both partial and full discrepancy (amber and red). Observe in all cases considered here, that the PRS indicates discrepancy (amber and red) more frequently than the PSI(YN).